OpenStudy (anonymous):

slope of the line passing through the point (-1, -3) and having an x-intercept of 2

OpenStudy (rock_mit182):

hypothesis y = x - 2 when x equals to 2: we have (2,0) the x-intercept now: P(-1,-3) y-(-1) =m(x-(-3)) m = 1 because of our hypothesis,

OpenStudy (rock_mit182):

y =x + 2 lol i guess i did someting wrong . .

OpenStudy (rock_mit182):

@shereenkhan can you solve this ?

OpenStudy (anonymous):

i am trying but i am not getting exact answer.

OpenStudy (rock_mit182):

so am I

OpenStudy (rock_mit182):

is a trick question

OpenStudy (anonymous):

yes u can :)

OpenStudy (anonymous):

slope is (-3-0)/(-1-2)=-3/-3=1

OpenStudy (anonymous):

any doubt?

OpenStudy (anonymous):

i am getting -1

OpenStudy (anonymous):

if A (x1,y1),B(x2,y2) be two points then slope of AB is (y1-y2)/(x1-x2)

OpenStudy (anonymous):

here (x1,y1)=(-1,-3) , (x2,y2)=(2,0)

OpenStudy (anonymous):

now you can easily understand

OpenStudy (anonymous):

yes i got it.

OpenStudy (anonymous):

thank U @Arr :)

OpenStudy (anonymous):

you are welcom